OpenStudy (midhun.madhu1987):

A father can do a job as fast as his 2 sons working together. If one son does the job in 3 hours and the other in 6 hours, how many hours does it take the father to do the job?

OpenStudy (midhun.madhu1987):

I am in need of the steps to find it...

OpenStudy (shinalcantara):

let 't' be the time needed for the job to be completed since you are given the rates of the two sons as: son1:1/3 son2:1/6 then the time for the father to finish the job would be: \[\frac{ 1 }{ 3 }t + \frac{ 1 }{ 6 }t = 1\] \[\frac{ 2t +t }{ 6 } = 1\] \[\frac{ 3t }{ 6 } = 1\] \[t = 2hours\]

OpenStudy (midhun.madhu1987):

Thanks a lot for ur prompt reply... :)